OpenStudy (anonymous):

There is a sales tax of $14 on an item that costs $147 before tax. A second item costs $84 before tax. What is the sales tax on the second item?

OpenStudy (anonymous):

@cassidyohrt

OpenStudy (anonymous):

@cassidyohrt

OpenStudy (studygurl14):

\[\frac{ 14 }{ 147 } = \frac{ T }{ 84 }\] \[14(84) = 147T\] \[1176 = 147T\] \[8 = T\] Tax = $8

OpenStudy (anonymous):

147 + 147*x = 147+14

OpenStudy (anonymous):

so its 8

OpenStudy (anonymous):

@cassidyohrt is that right

OpenStudy (anonymous):

The sales tax will be the same percent rate for both items. For the 1st item, then 14/147=0.095 0.095*84=8 then answer is 8

OpenStudy (studygurl14):

yes 8 is the answer @kaylee.carmestro